OpenStudy (anonymous):

5(2x + 1) = 3(5x - 4)

OpenStudy (anonymous):

10x + 5 = 15x - 12 Then, x = 17/5

OpenStudy (anonymous):

multiply 5 to 2x and 1 and 3 to 5x and -4 to get 1Ox +5=15x-12 5x=17 x=17/5